For the matrix A, find (if possible) a nonsingular matrix P such that p-1AP is diagonal. (If not possible, enter IMPOSSIBLE.)PLZ SOLVE BOTH WRONG PARTS
Answer #1

# A= [?] The characteristic. polynomial are 2,1- 92-92 +8 -8 = 0 → 4 (1-3) = 0 8-7 - 2 - 4 1-7 X X a = 0,9The eigen natues 1,=0, A2=9 a=0 8 [:] [3] [:)] 82 – 29 ,-4x+y = 0 If we take a=1, then y=4 The rector 1 4 eigen rector 9 - 9IPIS 1 +8=9 Adj P: 14-21 fa ) 5 (41) . Adi P PL 219 19 19 -4/9 yo 2/9 -4/9 29 PAP - - -690)6 ) - (83) 11 9 Cans) 2 # A= - 2> (2-2) (9-4) (9-1)=0 :.2= 1, 2,4 The eigen values 2,= 1,92-2, 23 = 4 2=1 -2 3 X ܘܘ ܘ o Y 0 2 -2 0 - 1 1, .. 2-29+37= 0, 24-2The eigen vector 7/2 -2 The eigen value a =1 then corresponding eigen vector 1 The eigen value a=2 then corresponding eigen v- o T- I P = adj 0 E 3 lo-1373/ -1 - 2 - 3 -% -3/ -1 -2 -3 -972 3/2 1PL 1 -1 2 3/2 - 1/2 o - 1/3 13 o 1/3 2/3 1-1 1 7/2 3/ -